Lying along the y-axis is a long, rigid wire which carries an 8.0 A current flowing in the positive direction.



(A) A 0.20 T constant magnetic field is directed along positive x-axis, find the magnetic force per unit length on the wire.



(B) If a 0.20 T constant magnetic force is directed 45 degrees from the +x-axis towards the +y-axis, solve for the magnetic force per unit length of the wire.

1
Expert's answer
2022-03-18T09:16:29-0400

(A) The magnetic force per unit length on the wire

"\frac{F}{l}=BI\sin\theta=0.20*8.0*\sin 90^\circ=1.6\:\rm N"

The force is directed toward negative z-axis.

(B) The magnetic force per unit length on the wire

"\frac{F}{l}=BI\sin\theta=0.20*8.0*\sin 45^\circ=1.1\:\rm N"

The force is directed toward negative z-axis.